# FIPS 204-Compatible Threshold ML-DSA via Masked Lagrange Reconstruction

### v3

@@ Abstract
[…]
Our technique overcomes the barrier that Lagrange coefficients grow as $\Theta(q)$ for moderate $T$, making individual contributions too large for rejection sampling. We introduce pairwise-canceling masks that hide these coefficients while preserving correctness. Unlike ECDSA threshold schemes, ML-DSA presents three additional challenges: the $\|\mathbf{z}\|_{\infty}$ check must pass after masking, the r0-check must not leak $c\mathbf{s}_{2}$, and the resulting Irwin-Hall nonce distribution must preserve EUF-CMA security. We solve all three with complete UC proofs.
We instantiate this technique in three deployment profiles. Profile P1 uses a TEE coordinator (3 rounds, 6ms). Profile P2 eliminates hardware trust via MPC (5 rounds, 22ms for small thresholds, up to 194ms for $T=16$). Profile P3+ uses lightweight 2PC with semi-asynchronous signing where signers precompute nonces offline and respond within a time window (2 logical rounds, 22ms), similar to FROST.
Our Rust implementation scales from 2-of-3 to 32-of-45 thresholds. P1 and P3+ achieve sub-100ms latency for all configurations; P2 trades higher latency for eliminating hardware trust. Success rates of $\approx$21–45% are comparable to single-signer ML-DSA’s $\approx$20–25%.
@@ 1.1 Motivation: The Threshold Gap
This leaves a gap: for $T\in[9,32]$, there is no known scheme that simultaneously achieves standard signature size, practical success rates, and FIPS 204 compatibility. This range is precisely where many real-world applications fall—organizations typically require 16-of-32 or similar thresholds for robust multi-party control.
@@ 1.2 Our Contribution
We present Masked Threshold ML-DSA, a threshold signature scheme for ML-DSA that supports arbitrary thresholds $T\leq N$ while producing standard-size signatures ($\approx 3.3$ KB for ML-DSA-65) identical to single-signer ML-DSA. Our scheme requires a two-honest requirement: signing sets must have size $|S|\geq T+1$ to ensure at least two honest parties for mask privacy. This addresses the threshold gap for $T\in[9,32]$ while remaining practical; configurations like $(3,5)$, $(11,15)$, and $(16,21)$ naturally satisfy $N\geq T+1$.
@@
The scheme achieves practical success rates comparable to single-signer ML-DSA (theoretical baseline $\approx 20\%$, observed $\approx$21–45% due to the more concentrated Irwin-Hall nonce distribution) and is fully verifier-compatible with FIPS 204: produced signatures are accepted by unmodified FIPS 204 verifiers. While our signer-side protocol is interactive and induces an Irwin-Hall nonce distribution rather than uniform, this affects only the signing process, not verification.
Prior approaches attempt to ensure each party’s contribution is individually small, but this hits a barrier: Lagrange coefficients $\lambda_{i}\approx 2^{T}$ grow exponentially, reaching $\Theta(q)$ for $T\geq 23$ in ML-DSA. If party $i$ naively computes $\lambda_{i}\cdot\mathbf{z}_{i}$ where $\|\mathbf{z}_{i}\|_{\infty}=O(\gamma_{1})$, the result exceeds the rejection threshold for any $\lambda_{i}>1$.
Our approach instead ensures that only the sum needs to be small, not individual contributions. While pairwise-canceling masks exist for ECDSA [CGGMP20], directly applying them to lattice signatures does not work because: (1) ML-DSA requires rejection sampling on $\|\mathbf{z}\|_{\infty}$, not just correctness; (2) the r0-check leaks $c\mathbf{s}_{2}$ which enables key recovery; (3) the Irwin-Hall nonce distribution must preserve EUF-CMA security. Our masked Lagrange reconstruction addresses these challenges:
  1. 1.
    Each pair of parties $(i,j)$ establishes a shared seed $\mathsf{seed}_{i,j}$ (e.g., via ML-KEM).
  2. 2.
    During signing, party $i$ computes a mask:
    $\mathbf{m}_{i}=\sum_{j>i}\mathsf{PRF}(\mathsf{seed}_{i,j},\mathsf{nonce})-\sum_{j<i}\mathsf{PRF}(\mathsf{seed}_{j,i},\mathsf{nonce})$
  3. 3.
    By construction, $\sum_{i\in S}\mathbf{m}_{i}=\mathbf{0}$ for any signing set SS—the masks pairwise cancel.
  4. 4.
    Party $i$ sends $\mathbf{U}_{i}=\mathbf{y}_{i}+\lambda_{i}\cdot c\cdot\mathbf{s}_{1,i}+\mathbf{m}_{i}$ instead of the unmasked value.
While each $\mathbf{U}_{i}$ may have arbitrary norm in [0,q1][0,q-1], their sum is:
$\sum_{i\in S}\mathbf{U}_{i}=\sum_{i}\mathbf{y}_{i}+c\cdot\sum_{i}\lambda_{i}\mathbf{s}_{1,i}+\sum_{i}\mathbf{m}_{i}=\mathbf{y}+c\cdot\mathbf{s}_{1}+\mathbf{0}=\mathbf{z}$
which correctly reconstructs the signature vector $\mathbf{z}$ and passes rejection sampling with probability comparable to single-signer ML-DSA (theoretical $\approx 20\%$, observed $\approx$21–45% for ML-DSA-65). We note that while the aggregated nonce $\mathbf{y}=\sum_{i}\mathbf{y}_{i}$ follows an Irwin-Hall distribution rather than uniform (see Definition Q.5), this does not affect EUF-CMA security and only introduces a quantifiable statistical difference that we analyze in Section 3.
@@
We prove security in the random oracle model. Unforgeability reduces to the Module-SIS problem, matching the security of single-signer ML-DSA, while privacy against coalitions of up to T1T-1 corrupted parties reduces to Module-LWE and PRF security.
@@
A TEE/HSM coordinator performs the r0-check and hint generation inside a trusted enclave, achieving optimal 3-round signing. We prove EUF-CMA security under Module-SIS and privacy under Module-LWE (Theorem 4.2).
@@ 1.3 Technical Overview
A trusted dealer creates Shamir shares of $\mathsf{sk}=(\mathbf{s}_{1},\mathbf{s}_{2})$ and establishes pairwise seeds via ML-KEM. The 3-round signing protocol has parties (1) commit to nonces $\mathbf{y}_{i}\in[-\gamma_{1}/|S|,\gamma_{1}/|S|]$, (2) exchange masked commitments to derive challenge $c$, and (3) send masked responses $\mathbf{U}_{i}=\mathbf{y}_{i}+\lambda_{i}\cdot c\cdot\mathbf{s}_{1,i}+\mathbf{m}_{i}$. The combiner aggregates $\mathbf{z}=\sum_{i}\mathbf{U}_{i}=\mathbf{y}+c\mathbf{s}_{1}$ (masks cancel) and outputs $\sigma=(\tilde{c},\mathbf{z},\mathbf{h})$ if $\|\mathbf{z}\|_{\infty}<\gamma_{1}-\beta$.
@@ 1.4 Comparison with Prior and Concurrent Work
Scheme $T$ Size Rnd FIPS Dishon. UC Async
del Pino-Niot [dPN25] $\leq 8$ 3.3 KB 3 $\times$ $\times$
Noise Flooding [BCDG21] $\infty$ 17 KB 3 $\times$ $\times$ $\times$
Ringtail [BMMSZ25] $\infty$ 13 KB 2 $\times$ $\times$ $\times$
@@
Bienstock et al. [BDEPT25] $\infty$ 3.3 KB 37–136 $\times$ $\times$
Celi et al. [CDENP26] $\leq 6$ 3.3 KB 6 $\times$ $\times$
Trilithium [Trilithium25] 2 3.3 KB 3 $\times$ $\times$
This work (P1) $\infty$ 3.3 KB 3 (TEE) $\times$
This work (P2) $\infty$ 3.3 KB 5 $\times$
This work (P3+) $\infty$ 3.3 KB 2 (1/2 CP)
Table 1: Comparison of threshold ML-DSA schemes. ✓= supported, $\times$ = not supported. “Dishon.” = dishonest-majority security (tolerates n1n-1 corruptions); Bienstock et al. require honest majority. “(TEE)” and “(1/2 CP)” indicate additional trust assumptions beyond threshold. “Async” = semi-asynchronous signer participation. P3+ requires 1 signer step + 1 server round = 2 logical rounds. P1 UC security assumes ideal TEE (Theorem 4.10).
@@
We contribute: (1) Masked Lagrange reconstruction, adapting pairwise-canceling masks [CGGMP20] to lattices while addressing rejection sampling, r0-check key leakage, and Irwin-Hall nonce security; (2) Three deployment profiles (P1-TEE, P2-MPC, P3+-2PC) with different trust/performance trade-offs, including the novel semi-async model for P3+; (3) Security proofs reducing to Module-SIS/LWE, including complete UC security proofs for P2 and P3+; (4) A reference implementation with benchmarks from $(2,3)$ to $(32,45)$ threshold configurations.
@@ 2.4 Shamir Secret Sharing
Lagrange coefficients $\lambda_{i}=(-1)^{i-1}\binom{T}{i}$ grow as $\approx 2^{T}/\sqrt{T}$. For $T\leq 8$, coefficients remain small ($\binom{8}{4}=70$), enabling short-coefficient schemes [dPN25]. For $T\geq 24$, coefficients exceed $q\approx 2^{23}$, aligning with the Ball-Çakan-Malkin bound [BCM21]. The obstacle is that $\lambda_{i}\cdot\mathbf{z}_{i}$ has norm $O(q\cdot\gamma_{1})\gg\gamma_{1}-\beta$, causing rejection.
@@ A $(T,N)$-threshold signature scheme consists of algorithms:
  • $\mathsf{Sign}(\{\mathsf{sk}_{i}\}_{i\in S},\mu)\to\sigma$: Interactive protocol among parties in $S$ (where $|S|\geq T+1$; see Section 4 for the two-honest requirement) to sign message $\mu$
@@
Remark 2.5 ($T+1$ Signing Requirement).
Our protocol requires $|S|\geq T+1$ signers (not $T$) to ensure at least two honest parties for mask privacy. This is standard for pairwise-mask schemes [CGGMP20, DKOSS24]; see Remark 4.1.
@@ 3.1 Overview
Our solution is based on two key observations:
@@
    Masking the coefficients: We add pairwise-canceling masks $\mathbf{m}_{i}$ to each party’s contribution. While individual contributions $\mathbf{U}_{i}$ may have arbitrary magnitude, the masks cancel in the sum, leaving the correct short result.
@@
    Separating nonce and secret: The nonce $\mathbf{y}_{i}$ and secret share $\mathbf{s}_{1,i}$ must be treated differently. Nonces are summed directly ($\mathbf{y}=\sum_{i}\mathbf{y}_{i}$), while secret shares are reconstructed via Lagrange interpolation ($\mathbf{s}_{1}=\sum_{i}\lambda_{i}\mathbf{s}_{1,i}$).
@@ 3.2 Threat Model and Deployment Profiles
Signing set size $|S|\geq T+1$ (two-honest requirement)
@@
Privacy goal T1T-1 privacy under Module-LWE

### v4

@@ Abstract
We present the first threshold ML-DSA (FIPS 204) scheme achieving statistical share privacy (no computational assumptions) with arbitrary thresholds while producing standard 3.3 KB signatures verifiable by unmodified implementations. Our primary technique is Shamir nonce DKG: parties jointly generate the signing nonce via a distributed key generation protocol, so that both the nonce and the long-term secret are degree-(T1)(T\!-\!1) Shamir sharings. The honest party’s response has conditional min-entropy exceeding $5\times$ the secret key entropy, requiring no computational assumptions. In coordinator-based profiles (P1, P3+), this eliminates the two-honest requirement for response privacy ($|S|\geq T$ suffices); in the fully distributed profile (P2), the existing mask-hiding condition $|S\setminus C|\geq 2$ applies. As a secondary technique, we introduce pairwise-canceling masks ($|S\setminus C|\geq 2$) for commitment and r0-check values, addressing the three challenges unique to lattice-based threshold signing: the $\|\mathbf{z}\|_{\infty}$ rejection sampling check, secure evaluation of the r0-check without leaking $c\mathbf{s}_{2}$, and EUF-CMA security under the resulting Irwin-Hall nonce distribution.
We instantiate these techniques in three deployment profiles with complete UC proofs. Profile P1 uses a TEE coordinator (3 online rounds plus 1 offline preprocessing round, 6ms). Profile P2 eliminates hardware trust via MPC (5 rounds, 22ms for small thresholds). Profile P3+ uses lightweight 2PC with semi-asynchronous signing (2 logical rounds, 22ms), where signers precompute nonces offline and respond within a time window. Our Rust implementation scales from 2-of-3 to 32-of-45 thresholds with sub-100ms latency (P1 and P3+) and success rates of $\approx$21–45%, comparable to single-signer ML-DSA.
@@ 1.1 Motivation: The Threshold Gap
This leaves a gap: for $T\in[9,32]$, there is no known scheme that simultaneously achieves standard signature size, practical success rates, and FIPS 204 compatibility. This range is precisely where many real-world applications fall (organizations typically require 16-of-32 or similar thresholds for robust multi-party control).
@@ 1.2 Our Contribution
We present Masked Threshold ML-DSA, the first threshold signature scheme for ML-DSA that achieves statistical share privacy (no computational assumptions) with arbitrary thresholds $T\leq N$ while producing standard-size signatures ($\approx 3.3$ KB for ML-DSA-65) identical to single-signer ML-DSA. Unlike prior pairwise-mask schemes that require a two-honest assumption ($|S|\geq T+1$), our protocol achieves response privacy with signing sets of size $|S|\geq T$ (the information-theoretic minimum for Shamir reconstruction) in coordinator-based profiles (P1, P3+), where individual masked commitments $\mathbf{W}_{i}$ are sent only to the coordinator rather than broadcast (Remark 3.7); commitment and r0-check mask hiding retains $|S\setminus C|\geq 2$ (Lemma 4.9), which applies to P2’s broadcast model. This addresses the threshold gap for $T\in[9,32]$ while remaining practical for configurations such as $(3,5)$, $(11,15)$, and $(16,21)$.
@@
The scheme achieves practical success rates comparable to single-signer ML-DSA (theoretical baseline $\approx 20\%$, observed $\approx$21–45% due to the more concentrated Irwin-Hall nonce distribution) and is fully verifier-compatible with FIPS 204. The aggregated nonce follows an Irwin-Hall distribution rather than uniform, but this does not affect EUF-CMA security and only introduces a quantifiable statistical difference analyzed in Section 3.
Our approach is based on two techniques. The primary technique is Shamir nonce DKG: parties jointly generate the signing nonce via a distributed key generation protocol, producing a degree-(T1)(T\!-\!1) Shamir sharing of $\mathbf{y}$ that matches the structure of the long-term secret $\mathbf{s}_{1}$. Each party then computes $\mathbf{z}_{i}=\mathbf{y}_{i}+c\cdot\mathbf{s}_{1,i}$ without applying the Lagrange coefficient, and the combiner reconstructs $\mathbf{z}=\sum_{i}\lambda_{i}\mathbf{z}_{i}=\mathbf{y}+c\mathbf{s}_{1}$. Because the adversary observes only T1T-1 evaluations of the honest party’s degree-(T1)(T\!-\!1) polynomial, the remaining degree of freedom provides an approximate one-time pad: the honest party’s response $\mathbf{z}_{h}$ has conditional min-entropy exceeding $5\times$ the secret key entropy (for $|S|\leq 17$), requiring no computational assumptions.
The secondary technique is pairwise-canceling masks, adapted from ECDSA [CGGMP20] to the lattice setting. These masks are used for the commitment values $\mathbf{W}_{i}$ and the r0-check shares $\mathbf{V}_{i}$ (which contain $\lambda_{i}\cdot c\cdot\mathbf{s}_{2,i}$), hiding individual contributions while preserving correctness in the sum. Unlike ECDSA, ML-DSA presents additional challenges: the r0-check leaks $c\mathbf{s}_{2}$ which enables key recovery, and the Irwin-Hall nonce distribution must preserve EUF-CMA security. We address both with complete UC proofs.
@@
We prove security in the random oracle model. Unforgeability reduces to the Module-SIS problem, matching the security of single-signer ML-DSA. Share privacy is statistical (no computational assumptions) via the Shamir nonce DKG, with conditional min-entropy exceeding $5\times$ the secret key entropy for $|S|\leq 17$; no computational assumptions are needed for response privacy. Commitment and r0-check privacy reduce to PRF security.
@@
A TEE/HSM coordinator performs the r0-check and hint generation inside a trusted enclave, achieving optimal 3-round signing. We prove EUF-CMA security under Module-SIS (Theorem 4.2), with statistical response privacy (Theorem 4.4) and PRF-based mask privacy (Lemma 4.9).
@@ 1.3 Technical Overview
A trusted dealer creates Shamir shares of $\mathsf{sk}=(\mathbf{s}_{1},\mathbf{s}_{2})$ and establishes pairwise seeds via ML-KEM. Before each signing attempt, parties execute a Shamir nonce DKG to jointly generate degree-(T1)(T\!-\!1) shares of the nonce $\mathbf{y}$ (this can be preprocessed offline). The 3-round online signing protocol then has parties (1) commit to their DKG nonce shares, (2) exchange masked commitments to derive challenge $c$, and (3) send responses $\mathbf{z}_{i}=\mathbf{y}_{i}+c\cdot\mathbf{s}_{1,i}$ (no Lagrange coefficient, no mask). The combiner applies Lagrange coefficients during aggregation: $\mathbf{z}=\sum_{i}\lambda_{i}\mathbf{z}_{i}=\mathbf{y}+c\mathbf{s}_{1}$, and outputs $\sigma=(\tilde{c},\mathbf{z},\mathbf{h})$ if $\|\mathbf{z}\|_{\infty}<\gamma_{1}-\beta$.
@@ 1.4 Comparison with Prior and Concurrent Work
Scheme $T$ Size Rnd FIPS Dishon. UC Async Privacy
del Pino-Niot [dPN25] $\leq 8$ 3.3 KB 3 $\times$ $\times$ Comp.
Noise Flooding [BCDG21] $\infty$ 17 KB 3 $\times$ $\times$ $\times$ Comp.
Ringtail [BMMSZ25] $\infty$ 13 KB 2 $\times$ $\times$ $\times$ Comp.
@@
Bienstock et al. [BDEPT25] $\infty$ 3.3 KB 37–136 $\times$ $\times$ Comp.
Celi et al. [CDENP26] $\leq 6$ 3.3 KB 6 $\times$ $\times$ Comp.
Trilithium [Trilithium25] 2 3.3 KB 3 $\times$ $\times$ Comp.
This work (P1) $\infty$ 3.3 KB 3 (TEE) $\times$ Stat.
This work (P2) $\infty$ 3.3 KB 5 $\times$ Stat.
This work (P3+) $\infty$ 3.3 KB 2 (1/2 CP) Stat.
Table 1: Comparison of threshold ML-DSA schemes. ✓= supported, $\times$ = not supported. “Dishon.” = dishonest-majority security (tolerates N1N-1 corruptions); Bienstock et al. require honest majority. “Privacy”: Comp. = computational (PRF-based), Stat. = statistical (no computational assumptions, min-entropy $\geq 5\times$ secret key entropy) for response shares $\mathbf{z}_{i}$; commitment/r0-check privacy uses PRF-based masks ($|S\setminus C|\geq 2$). “(TEE)” = dishonest-majority via TEE trust (not purely cryptographic); “(1/2 CP)” = via 1-of-2 computation-party honest assumption. P3+ = 1 signer step + 1 server round. P1 UC assumes ideal TEE (Theorem 4.13). Plus 1 offline preprocessing round (nonce DKG).
@@
We contribute: (1) Shamir nonce DKG for threshold ML-DSA, achieving statistical response privacy (no computational assumptions) and in coordinator-based profiles (P1, P3+), eliminating the two-honest requirement for response shares ($|S|\geq T$ suffices; see Remark 3.7)—to our knowledge, the first threshold lattice-based signature with statistical share privacy and $|S|=T$ signing sets; (2) Pairwise-canceling masks adapted to the lattice setting for commitment and r0-check values, addressing rejection sampling and r0-check key leakage; (3) Three deployment profiles (P1-TEE, P2-MPC, P3+-2PC) with different trust/performance trade-offs, including the novel semi-async model for P3+; (4) Complete security proofs, with unforgeability reducing to Module-SIS, statistical privacy via Shamir nonce DKG (Theorem 4.4), and UC security proofs for all profiles; (5) A reference implementation with benchmarks from $(2,3)$ to $(32,45)$ threshold configurations.
@@ 2.3 ML-DSA (FIPS 204)
  • $\omega=55$: hint weight bound (max number of $1$-entries in hint $\mathbf{h}$)
  • $d=13$: dropping bits in $\mathsf{Power2Round}$ ($\mathbf{t}=\mathbf{t}_{1}\cdot 2^{d}+\mathbf{t}_{0}$)
@@ 2.4 Shamir Secret Sharing
For the standard evaluation points $\{1,2,\ldots,T\}$, Lagrange coefficients satisfy $\lambda_{i}=(-1)^{i-1}\binom{T}{i}$, growing as $\approx 2^{T}/\sqrt{T}$. For $T\leq 8$, coefficients remain small ($\binom{8}{4}=70$), enabling short-coefficient schemes [dPN25]. For $T\geq 24$, coefficients exceed $q\approx 2^{23}$, aligning with the Ball-Çakan-Malkin bound [BCM21]. The obstacle is that $\lambda_{i}\cdot\mathbf{z}_{i}$ has norm $O(q\cdot\gamma_{1})\gg\gamma_{1}-\beta$, causing rejection.
@@ A $(T,N)$-threshold signature scheme consists of algorithms:
  • $\mathsf{Sign}(\{\mathsf{sk}_{i}\}_{i\in S},\mu)\to\sigma$: Interactive protocol among parties in $S$ (where $|S|\geq T$) to sign message $\mu$
@@
Remark 2.5 (Signing Set Size).
With Shamir nonce DKG and coordinator-based communication (Remark 3.7), Profiles P1 and P3+ require only $|S|\geq T$ signers—the information-theoretic minimum for Shamir reconstruction. This improves upon standard pairwise-mask schemes [CGGMP20, DKOSS24] that require $|S|\geq T+1$. The nonce DKG provides statistical share privacy (no computational assumptions) even with a single honest party (see Theorem 4.4 and Remark 4.1). In P2’s broadcast model, $|S\setminus C|\geq 2$ is required for commitment privacy (Lemma 4.9).
@@ 3.1 Overview
Our solution is based on two key techniques:
@@
    Shamir nonce DKG: Parties jointly generate the nonce $\mathbf{y}$ via a Shamir-style distributed key generation, so that both $\mathbf{y}$ and the long-term secret $\mathbf{s}_{1}$ are degree-(T1)(T\!-\!1) Shamir sharings. Each party computes $\mathbf{z}_{i}=\mathbf{y}_{i}+c\cdot\mathbf{s}_{1,i}$ (without Lagrange coefficients), and the combiner reconstructs $\mathbf{z}=\sum_{i}\lambda_{i}\mathbf{z}_{i}=\mathbf{y}+c\mathbf{s}_{1}$. This achieves statistical share privacy (conditional min-entropy $\geq 5\times$ secret key entropy for $|S|\leq 17$, requiring no computational assumptions), in coordinator-based profiles (P1, P3+), eliminates the two-honest requirement ($|S|\geq T$ suffices; see Remark 3.7), and requires no pairwise masks on the response $\mathbf{z}_{i}$.
@@
    Pairwise-canceling masks: For the commitment values $\mathbf{W}_{i}$ and the r0-check shares $\mathbf{V}_{i}$, we add pairwise-canceling masks that hide individual contributions while preserving correctness in the sum.
@@ 3.2 Threat Model and Deployment Profiles
Signing set size $|S|\geq T$
@@
Privacy goal Statistical share privacy (Theorem 4.4)
@@ 3.3 Setup Phase
3.4 Shamir Nonce DKG
The primary technique enabling statistical share privacy is the Shamir nonce DKG: parties jointly generate the signing nonce $\mathbf{y}$ as a degree-(T1)(T\!-\!1) Shamir sharing, matching the structure of the long-term secret $\mathbf{s}_{1}$. This is executed once per signing attempt and can be preprocessed offline.
Algorithm 2 $\mathsf{NonceDKG}(S)$: Distributed Nonce Generation
1:Input: Signing set $S$ with $|S|\geq T$
2:for each party $i\in S$ in parallel do
3:  Sample secret contribution $\hat{\mathbf{y}}_{i}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}\{-\lfloor\gamma_{1}/|S|\rfloor,\ldots,\lfloor\gamma_{1}/|S|\rfloor\}^{n\ell}$
4:  Construct degree-(T1)(T\!-\!1) polynomial over $R_{q}^{\ell}$:
5:  $\mathbf{f}_{i}(x)=\hat{\mathbf{y}}_{i}+\mathbf{a}_{i,1}x+\cdots+\mathbf{a}_{i,T-1}x^{T-1}$
6:  where $\mathbf{a}_{i,k}\stackrel{{\scriptstyle\mathdollar}}{{\leftarrow}}R_{q}^{\ell}$ for $k=1,\ldots,T\!-\!1$
7:  Send $\mathbf{f}_{i}(j)$ to each party $j\in S$ over a private channel $\triangleright$ Encrypted via pairwise ML-KEM seeds
8:end for
9:for each party $j\in S$ do
10:  Compute nonce share: $\mathbf{y}_{j}=\sum_{i\in S}\mathbf{f}_{i}(j)$
11:end for
The combined polynomial is $\mathbf{F}(x)=\sum_{i\in S}\mathbf{f}_{i}(x)$, with $\mathbf{F}(0)=\sum_{i}\hat{\mathbf{y}}_{i}=\mathbf{y}$ (the nonce) and $\mathbf{F}(j)=\mathbf{y}_{j}$ (the shares). This is a valid degree-(T1)(T\!-\!1) Shamir sharing of $\mathbf{y}$ over $R_{q}^{\ell}$.
Remark 3.1 (Why Uniform Higher-Degree Coefficients).
The coefficients $\mathbf{a}_{i,k}$ for $k\geq 1$ are sampled uniformly over $R_{q}^{\ell}$ (not from a short distribution). This is essential: the adversary observes T1T-1 evaluations of $\mathbf{f}_{h}$ (the honest party’s polynomial), leaving exactly one degree of freedom. The uniform prior distribution of $\mathbf{a}_{h,T-1}$ ensures high entropy in this free parameter; combined with the bounded-nonce constraint, the honest party’s share $\mathbf{f}_{h}(x_{h})$ retains conditional min-entropy $\geq n\ell\cdot\log_{2}(2\gamma_{1}/|S|)$ bits per session. See Theorem 4.4 and Remark 4.6.
Remark 3.2 (Communication Cost).
The nonce DKG requires $|S|\cdot(|S|-1)$ point-to-point messages, each of size $n\ell\cdot\lceil\log_{2}q\rceil$ bits ($\approx 3.8$ KB for ML-DSA-65). For a $(3,5)$-threshold with $|S|=3$, this is 6 messages totaling $\approx 23$ KB. The DKG round is independent of the message to be signed and can be executed offline as a preprocessing step, so online signing latency is unchanged.